Allocated function: y = A cos kt

Values assigned: A = 4, k = 5

Graph of displacement y (metres) against time t (seconds) for time interval from t = 0s to t = 2π/k seconds.

Q: Calculate the turning points of the function using differential calculus and show for each turning point whether it is a maximum, a minimum or a point of inflexion, by using the second derivative of the function.


1
Expert's answer
2021-05-12T03:50:14-0400

"y=A\cos(kt)"

but A=4 and k=5

"\implies y=4 \cos(5t)"

at turning points,first derivative is zero

"\implies {dy\over dx}=4(-\sin(5t)*5)=-20 \sin(5t)=0"

"\implies \sin(5t)=0" ......................(i)

let "5t=\theta \implies \sin(\theta)=0"

from the unit circle,we have two angles whose sine is "\theta" ,that is 0 and "\pi"

"\implies \theta=0+2\pi c=2\pi c"

"\theta=\pi +2 \pi c" , where c in an interger

but "\theta=5t"

"\theta=0+2 \pi c=2 \pi c \implies 2 \pi c=5t \implies t={2 \pi c \over 5}..............(ii)"

"\theta=\pi + 2 \pi c \implies \pi+2 \pi c=5t \implies t={\pi\over 5}+{2 \pi c\over 5}.................(iii)"

we then use equation (ii) and (iii) avove to find all of the infinitely many solutions for equation (i)

we are asked to find those solutions which are in the interval "[0 , {2 \pi\over 5}]"

this means that we shall stop at "t={2 \pi\over 5}"

we then substitute v arious values of c to equation (ii) and (iii) as follows

when "c=0"

"t={2 \pi (0)\over 5}=0" from equation (ii)

"t={\pi\over 5}+{2 \pi (0)\over 5}={\pi\over 5}" from equation (iii)

when "c=1"

"t={2 \pi (1)\over 5}={2 \pi\over 5}" from equation (ii)

"t={\pi\over 5}+{2 \pi (1)\over 5}={3\pi\over 5}" from equation (iii)

from interval "[0 , {2 \pi\over 5}]"

t=0 , "t={\pi\over 5}" , "t={2 \pi\over 5}"

we then find y values using the following equation

"y=4 \cos(5t)"

when "t=0"

"y=4 \cos(5*0)=4" "\implies" turning point "(0 , 4)"

when "t={\pi\over 5}"

"y=4 \cos(5*{\pi\over 5})=-4" "\implies" turning point "({\pi\over 5} ,- 4)"

when "t={2 \pi\over 5}"

"y=4 \cos(5*{2 \pi\over 5})=4" "\implies" turning point "({2\pi\over 5} , 4)"

we use the second derivative to determine the nature of the turning points

"{d^2y\over dt^2}=-20(\cos(5t)*5)"

"=-100\cos(5t)"

when "t=0"

"{d^2y\over dt^2}=-100\cos(5*0)=-100\lt0" "\implies (0 , 4)" is a local maximum

when "t={\pi\over 5}"

"{d^2y\over dt^2}=-100\cos(5*{\pi\over 5}=100\gt0" "\implies ({\pi\over 5} , -4)" is a local minimum

when "t={2 \pi\over 5}"

"{d^2y\over dt^2}=-100\cos(5*{2\pi\over 5})=-100\lt0" "\implies (0 , 4)" is a local maximum